OpenStudy (anonymous):

Please help!!! Write as a sum of partial fractions: x^(2)+1/x^(3)+x^(2)-5x+3

OpenStudy (comm.dan):

@jim_thompson5910

OpenStudy (anonymous):

I'm not sure what the final answers are? I turned it into a system of equations and got A=-1/2 B=-2/3 C=3/2

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

It said these were the answers....

OpenStudy (anonymous):

\[\frac{x^2+1}{x^3+x^2-5x+3}\]right ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

think your answer is not right

OpenStudy (anonymous):

My book has that answer but I have the answer you gave me.

OpenStudy (anonymous):

it is hard to tell when you only write A =, B = , C = , because it is not clear what the denominators are you need to specify what the denominators of those terms are

OpenStudy (anonymous):

ok good, because those answers are wrong, so there is either a typo in the book or in the question we can check it easily

OpenStudy (anonymous):

It had this as the answer (-1/2)/(x-1),(-2/3)/(x-1)^2,(3/2)/(x+3)

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

wolfram never lies sometimes it is hard to interpret what it says, but not in this case we can add those up and see that you do not get what you started with

OpenStudy (anonymous):

My book is so annoying!

OpenStudy (anonymous):

don't fret about it the problems were made up by underpaid overworked graduate students who do not check every last detail probably made up not even for your book, but for some question bank

OpenStudy (anonymous):

Yeah! I just kept thinking I was doing everything wrong. So I can check by hand to make sure I'm correct?

OpenStudy (anonymous):

sure add if you like, but personally i would not bother with all the algebra, plug it in to wolfram and check that way if wolfram says you are right, you are right, and if it says you are wrong... then you are wrong in this case it says the book is wrong and you are right

OpenStudy (anonymous):

I have a test tomorrow over it so I'm worried I don't know how to do it :/

OpenStudy (anonymous):

do it the way we did it yesterday, and don't fret that the book has the wrong answer

OpenStudy (anonymous):

trying to say don't give it any more thought, don't get annoyed, it is just a garden variety mistake in the book happens all the time put it aside and moveon.org